# Data-Driven Design Rules for TADF Emitters from a High-Throughput Screening of $747$ Molecules

### v1

@@ Abstract
I Introduction
@@
Organic Light-Emitting Diodes (OLEDs) based on Thermally Activated Delayed Fluorescence (TADF) have defined the third generation of electroluminescent materials, offering a viable path to achieving near-unity internal quantum efficiency (IQE) without resorting to the expensive and scarce noble metals found in phosphorescent emitters [1, 2, 3]. The TADF mechanism hinges on the efficient harvesting of non-emissive triplet excitons ($T_{1}$) by up-converting them to the emissive singlet state ($S_{1}$) through Reverse Intersystem Crossing (RISC). This process requires satisfying both a thermodynamic condition—a minimized singlet-triplet energy gap ($\Delta E_{\text{ST}}\qty$$\lesssim\qty{0.2}{}$)—and a kinetic condition: a sufficiently high rate of RISC ($k_{\text{RISC}}$) [4, 5].
@@
The rational design of high-performance TADF emitters is a formidable challenge, dominated by a fundamental trade-off. In conventional Donor-Acceptor (D-A) architectures, achieving a small $\Delta E_{\text{ST}}$necessitates a significant spatial separation of the frontier molecular orbitals to minimize the electron exchange energy [6]. However, this separation inherently reduces the orbital overlap, which not only suppresses the oscillator strength ($f$) and thus the radiative decay rate, but can also hinder the spin-orbit coupling (SOC) necessary for efficient RISC [7, 8]. Successfully navigating this delicate balance requires exploring a vast and complex chemical space [9].
Computational chemistry is an indispensable tool for navigating this landscape, yet it presents its own dilemma. High-fidelity methods are computationally prohibitive for screening the thousands of potential candidates required for data-driven discovery [10, 11, 12]. This scalability bottleneck has historically limited computational studies to small, homologous series of molecules, hindering the extraction of universal design principles.
@@
To bridge this gap, we employ a high-throughput virtual screening (HTVS) protocol based on accelerated semi-empirical methods, which we rigorously validated on a diverse dataset of $747$ experimentally known TADF emitters in our preceding companion paper (hereafter, Article 1) [13]. That study established our framework as a powerful and validated tool for large-scale molecular ranking and trend analysis.
In this article, we leverage our validated HTVS protocol to move beyond benchmarking and perform a large-scale, data-driven analysis aimed at elucidating the fundamental structure-property relationships that govern TADF performance. By analyzing the results from the entire 747-molecule dataset, we aim to distill the complex interplay of molecular structure, conformation, and electronic properties into a set of clear, quantitative, and actionable design guidelines. This study, therefore, seeks to: (i) establish a quantitative performance hierarchy among diverse molecular architectures (D-A, D-A-D, and MR systems); (ii) derive statistically robust design rules linking key structural parameters to both thermodynamic ($\Delta E_{\text{ST}}$) and kinetic (SOC, $k_{\text{RISC}}$) factors; (iii) identify and characterize distinct molecular families through data-driven clustering analysis; and (iv) propose a curated list of high-priority candidate molecules predicted to exhibit exceptional TADF properties. By translating a massive computational dataset into actionable design intelligence, this work provides a unified framework to guide and accelerate the discovery of next-generation TADF emitters.
@@
This paper is structured as follows: Section II details our methodological framework. The main results are presented in the subsequent sections, including an analysis of the architectural landscape, a classification of molecular families, and a deep dive into the MR-TADF design paradigm. Finally, Section IV summarizes our conclusions and proposes future directions for TADF emitter design.
II Methodological framework for large-scale analysis
@@
All structure-property relationships presented in this work were derived from a large-scale dataset of $747$ experimentally known TADF emitters, processed using a consistent and validated high-throughput virtual screening (HTVS) protocol. The full details, validation, and limitations of this computational framework are established in our preceding work (Article 1) [13]. Here, we provide a concise summary of the essential methodological components and their justifications.
II.1 Dataset and molecular structures
@@
The dataset comprises a diverse set of $747$ TADF molecules curated from the literature [14]. To facilitate systematic analysis, molecules were classified into distinct architectural families (e.g., D-A, D-A-D, Multi-Resonance) using automated SMARTS pattern matching, followed by manual verification.
@@
Initial 3D structures were generated from SMILES strings, and a rigorous conformational search was performed for each molecule using the CREST program coupled with the GFN2-xTB semi-empirical Hamiltonian [15, 16]. The lowest-energy conformer was then fully optimized at the GFN2-xTB level to obtain the final ground-state ($S_{0}$) geometry. This tight-binding method is specifically parameterized for accurate molecular structures and noncovalent interactions, making it well-suited for the complex and often flexible systems studied here [17].
II.2 Methodological approximations and justifications
II.2.1 Hybrid protocol and the vertical approximation compromise
@@
The HTVS protocol relies on a hybrid computational strategy: ground-state ($S_{0}$) geometries are optimized using GFN2-xTB [16, 18], while excited-state properties ($S_{1}$, $T_{1}$) are calculated using the simplified methods, sTDA-xTB/sTD-DFT-xTB [19, 20]. This methodology implicitly employs the vertical approximation.
@@
II.2.2 Limitations of the solvation model
The study of excited-state properties was performed using the ground-state-based ALPB implicit solvation model for toluene. We acknowledge that for the highly polarized charge-transfer (CT) states characteristic of TADF emitters, a state-specific solvation model (e.g., SS-PCM) is theoretically required for quantitative accuracy. However, such approaches are computationally prohibitive for HTS. The ALPB model is considered sufficient for screening as it has been shown that while the absolute solvation effect may be approximated, the relative ordering of molecules and the overall structure-property trends are conserved in the condensed phase.
II.2.3 Specific limitations for multi-resonance emitters
It is important to note that Multi-Resonance (MR) TADF emitters are inherently susceptible to multi-reference character [22], a feature that current single-reference based TD-DFT and its semi-empirical variants (sTDA-xTB/sTD-DFT-xTB) may not fully capture. Consequently, while our HTS trends provide excellent initial screening recommendations, definitive quantitative validation for MR-TADF candidates should rely on post-hoc applications of high-level methods such as SCS-CC2 or STEOM-DLPNO-CCSD [23].
II.3 Derivation of structure-property descriptors and theoretical framework
@@
To translate the large computational dataset into actionable design rules, donor-acceptor (D-A) torsional angles were automatically extracted from all optimized geometries. To quantify the degree of charge-transfer and its relation to geometry, we calculated the HOMO-LUMO overlap integral ($S^{\prime}_{\text{HL}}$) and the distance between the centroids of the hole and electron densities ($\Delta r_{\text{CT}}$) using the Multiwfn package [24, 25]. Principal Component Analysis (PCA) was performed on the standardized matrix of all descriptors to identify the underlying variables governing TADF performance.
@@
This analysis is grounded in the theoretical framework of TADF efficiency, which depends not only on a small $\Delta E_{\text{ST}}$(thermodynamics) but also on a high rate of reverse intersystem crossing ($k_{\text{RISC}}$), a kinetic parameter governed by spin-orbit coupling (SOC). According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation) [26, 27]. Our chosen descriptors—particularly the D-A torsional angle—are direct proxies for tuning this CT/LE character balance. While our protocol does not compute SOC explicitly due to its prohibitive cost, analyzing these structural and electronic descriptors provides a robust, qualitative assessment of the kinetic factors that, alongside $\Delta E_{\text{ST}}$, govern overall TADF performance.
@@
III Results and discussion: from statistical trends to rational design principles
@@
This section leverages the large-scale dataset of $747$ TADF emitters, processed with the validated HTVS protocol detailed in our companion paper (Article 1)[13], to extract quantitative, statistically robust structure-property relationships. Our analysis transitions from broad statistical trends to specific, actionable molecular design rules, with a focus on providing a deep mechanistic understanding that connects molecular structure to the key photophysical parameters governing TADF efficiency.
III.1 Low-dimensional nature of the TADF design space
@@
III.2 Architectural hierarchy defines TADF potential
@@
III.3 Data-driven clustering reveals distinct molecular families
To complement the structure-based classification, we performed an unsupervised K-means clustering based on the molecules’ photophysical properties. This data-driven approach identifies four distinct molecular families, as shown in Figure 3:
@@ III.3 Data-driven clustering reveals distinct molecular families
III.4 Multi-resonance TADF: a distinct paradigm for blue emitters
@@
However, as noted in the methodology, it is important to interpret these results with caution. The inherent electronic complexity of MR-TADF systems makes them prone to multi-reference character [22], which is not fully captured by the single-reference nature of the semi-empirical methods used in this HTS. The trends presented should therefore be interpreted as a valuable guide for compound ranking rather than a source of definitive quantitative predictions, with final validation requiring higher levels of theory.
@@
III.5 Deconstructing performance: geometric and electronic design levers
III.5.1 Torsional angle as a primary control parameter
@@
III.5.2 Quantifying the torsion-induced electronic trade-off
This geometric control is directly reflected in the electronic structure. As shown in Figure 5, twisting the molecule reduces the HOMO-LUMO overlap ($S^{\prime}_{\text{HL}}$), which correlates strongly with a smaller $\Delta E_{\text{ST}}$(Spearman $\rho=-0.45$). However, this comes at a cost: the oscillator strength, essential for a high radiative rate, is directly proportional to this overlap ($\rho=0.78$). This visualization makes the fundamental TADF design trade-off explicit: minimizing the energy gap via orbital separation inherently suppresses emission intensity. An optimal region for high-performance emitters lies where there is intermediate overlap (0.15 to 0.300.150.30) that balances these competing requirements.
To move beyond this qualitative picture and establish a more rigorous quantitative link, future work could employ methods such as Natural Transition Orbitals (NTOs) or the RespA procedure [29, 30]. These tools would allow for a compact representation of the virtual excitations and rigorously connect the molecular geometry to the precise charge-transfer character (e.g., via centroid distance $\Delta r$) and the ensuing coupling efficiency [31].
@@ III.5.2 Quantifying the torsion-induced electronic trade-off
III.6 Implications for reverse intersystem crossing ($k_{\text{RISC}}$)
While a small $\Delta E_{\text{ST}}$is the primary thermodynamic requirement for TADF, the overall efficiency is governed by the kinetics of RISC. While direct calculation of the $k_{\text{RISC}}$ rate is outside the scope of HTS, the derived structural principles must be explicitly connected to the kinetics of TADF. Within the Marcus theory framework, $k_{\text{RISC}}$ is proportional to the square of the spin-orbit coupling (SOC) matrix element and exponentially dependent on $\Delta E_{\text{ST}}$and the reorganization energy ($\lambda$) [32].
According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation, 1CT $\leftrightarrow$ 3LE) [26]. This creates a mechanistic challenge: the very act of creating a pure CT state in $S_{1}$ to minimize $\Delta E_{\text{ST}}$can lead to a $T_{1}$ state that also has pure CT character, making the transition spin-forbidden (1CT $\leftrightarrow$ 3CT) and suppressing $k_{\text{RISC}}$.
@@ III.6 Implications for reverse intersystem crossing ($k_{\text{RISC}}$)
HOMO-LUMO overlap Directly modulates $\Delta E_{\text{ST}}$ (thermodynamics) 0.15 to 0.300.150.30
@@
III.7 Synthesis of a data-driven design strategy
The combination of these large-scale statistical analyses allows us to formulate a set of quantitative, data-driven design guidelines for accelerating the discovery of novel TADF emitters, summarized in Table 3. These rules provide a clear roadmap for molecular engineering, from selecting a high-performance architecture to fine-tuning its properties via geometric and electronic modifications.
@@ III.7 Synthesis of a data-driven design strategy
HOMO-LUMO Overlap ($S^{\prime}_{\text{HL}}$) 0.15 to 0.300.150.30 Navigates the trade-off between $\Delta E_{\text{ST}}$  and oscillator strength.
@@
III.8 Future outlook and data exploitation
III.8.1 The HTS dataset as a machine learning substrate
@@
III.8.2 Actionable design targets
@@
IV Conclusions
In this work, we performed a large-scale computational screening of $747$ TADF emitters to extract quantitative, statistically robust structure-property relationships for guiding molecular design. By systematically analyzing the impact of molecular architecture, conformational geometry, and electronic structure, we have moved beyond anecdotal evidence to establish a data-driven framework that integrates both the thermodynamic and kinetic principles of TADF engineering.

### v2

# Data-Driven Design Guidelines for TADF Emitters from a High-Throughput Screening of $747$ Molecules

@@ Abstract
keywords:
TADF, OLED, high-throughput screening, structure-property relationships, singlet-triplet gap, xTB, molecular design
INTRODUCTION
@@
Organic Light-Emitting Diodes (OLEDs) based on Thermally Activated Delayed Fluorescence (TADF) have defined the third generation of electroluminescent materials, offering a viable path to achieving near-unity internal quantum efficiency (IQE) without resorting to the expensive and scarce noble metals found in phosphorescent emitters 1, 2, 3. The TADF mechanism hinges on the efficient harvesting of non-emissive triplet excitons ($T_{1}$) by up-converting them to the emissive singlet state ($S_{1}$) through Reverse Intersystem Crossing (RISC). This process requires satisfying both a thermodynamic condition—a minimized singlet-triplet energy gap ($\Delta E_{\text{ST}}\qty$$\lesssim\qty{0.2}{}$)—and a kinetic condition: a sufficiently high rate of RISC ($k_{\text{RISC}}$) 4, 5.
@@
The rational design of efficient TADF emitters faces an inherent trade-off that complicates optimization. In conventional Donor-Acceptor (D-A) architectures, achieving a small $\Delta E_{\text{ST}}$necessitates a significant spatial separation of the frontier molecular orbitals to minimize the electron exchange energy 6. However, this separation inherently reduces the orbital overlap, which not only suppresses the oscillator strength ($f$) and thus the radiative decay rate, but can also hinder the spin-orbit coupling (SOC) necessary for efficient RISC 7, 8. Optimizing both properties simultaneously demands exploration across diverse molecular scaffolds 9.
Computational methods can accelerate this search, but they introduce a cost-accuracy trade-off. High-fidelity methods are computationally prohibitive for screening the thousands of potential candidates required for data-driven discovery 10, 11, 12. This cost barrier has confined most studies to small molecular series. Universal design guidelines are therefore still lacking.
@@
Here, we employ a high-throughput virtual screening (HTVS) protocol based on accelerated semiempirical methods. We validated this approach on a data set of $747$ experimentally known TADF emitters in our companion paper (hereafter, Article 1 13), which established its reliability for large-scale molecular ranking and trend analysis.
In this article, we leverage our validated HTVS protocol to move beyond benchmarking and perform a large-scale analysis of structure-property relationships governing TADF performance. By analyzing the entire 747-molecule data set, we aim to establish quantitative design guidelines from the interplay of molecular structure, conformation, and electronic properties. We aim to establish an architecture ranking, derive structure-property rules for both $\Delta E_{\text{ST}}$ and $k_{\text{RISC}}$, classify molecules by clustering, and propose synthetic targets with favorable predicted properties. From this data set, we extract practical design guidelines. These can inform the synthesis of improved TADF materials.
@@
This paper is structured as follows: Section METHODOLOGICAL FRAMEWORK FOR LARGE-SCALE ANALYSIS details our methodological framework. The main results are presented in the subsequent sections, including an analysis of the architectural landscape, a classification of molecular families, and a deep dive into the MR-TADF design paradigm. Finally, Section CONCLUSIONS summarizes our conclusions and proposes future directions for TADF emitter design.
METHODOLOGICAL FRAMEWORK FOR LARGE-SCALE ANALYSIS
@@
All structure-property relationships presented in this work were derived from a large-scale data set of $747$ experimentally known TADF emitters, processed using a consistent and validated high-throughput virtual screening (HTVS) protocol. The full details, validation, and limitations of this computational framework are established in our preceding work (Article 1 13). Here, we provide a concise summary of the essential methodological components and their justifications.
Data Set and Molecular Structures
@@
The data set comprises a diverse set of $747$ TADF molecules curated from the literature 14. To facilitate systematic analysis, molecules were classified into distinct architectural families (e.g., D-A, D-A-D, Multiresonance) using automated SMARTS pattern matching, followed by manual verification.
@@
Initial 3D structures were generated from SMILES strings, and a rigorous conformational search was performed for each molecule using the CREST program coupled with the GFN2-xTB semiempirical Hamiltonian 15, 16. The lowest-energy conformer was then fully optimized at the GFN2-xTB level to obtain the final ground-state ($S_{0}$) geometry. This tight-binding method is parametrized for molecular structures and noncovalent interactions, which is appropriate for the flexible systems in our data set 17.
Methodological Approximations and Justifications
Hybrid Protocol and the Vertical Approximation Compromise
@@
The HTVS protocol relies on a hybrid computational strategy: ground-state ($S_{0}$) geometries are optimized using GFN2-xTB 16, 18, while excited-state properties ($S_{1}$, $T_{1}$) are calculated using the simplified methods, sTDA-xTB/sTD-DFT-xTB 19, 20. This methodology implicitly employs the vertical approximation.
@@
Solvation Model Selection and Trade-offs
The treatment of solvation effects in excited-state calculations presents a fundamental trade-off between computational efficiency and physical accuracy, particularly for charge-transfer (CT) states characteristic of TADF emitters. This study employed the ground-state-based GBSA implicit solvation model for toluene, which warrants detailed discussion of its limitations and the rationale for this choice in high-throughput screening applications.
Ground-state Vs. State-specific Solvation Models
Traditional implicit solvation models like GBSA utilize the ground-state electron density to compute solvation energies for all electronic states.22 While computationally efficient, this approximation becomes problematic for CT states where significant charge redistribution occurs upon excitation.23 The solvent response to electronic excitation can differ substantially from ground-state solvation, as the reorganized electron density in CT states creates different electrostatic interactions with the surrounding medium.24
State-specific solvation models, such as the state-specific polarizable continuum model (SS-PCM), address this limitation by computing separate solvation energies for each electronic state using the corresponding electron density.24, 25 For TADF emitters, where the $S_{1}$ and $T_{1}$ states often exhibit pronounced CT character with substantial charge separation between donor and acceptor moieties, SS-PCM can provide significantly more accurate solvation energies.6
Implications for Tadf Properties
The choice of solvation model has direct implications for predicted TADF properties. Solvent polarity modulates the singlet-triplet energy gap ($\Delta E_{\text{ST}}$) through differential solvation of the $S_{1}$ and $T_{1}$ states.26 CT states, with their spatially separated charges, typically exhibit stronger solvation effects than locally excited states. Consequently, ground-state-based models may systematically under- or overestimate $\Delta E_{\text{ST}}$  values, particularly in polar solvents.
Recent computational studies of TADF emitters using state-specific approaches have demonstrated that solvent effects can alter $\Delta E_{\text{ST}}$  by 0.1-0.3 eV compared to gas-phase calculations, with the magnitude depending on the degree of charge separation and solvent polarity.6 Ground-state-based models may capture the overall trend but with reduced quantitative accuracy.
Computational Cost Considerations
The primary limitation of state-specific solvation models is computational cost. SS-PCM requires separate self-consistent field calculations for each electronic state, typically increasing computational time by a factor of 2-3 compared to ground-state models.25 For high-throughput screening of 747 molecules with multiple conformers, this translates to thousands of additional calculations, making SS-PCM computationally prohibitive within reasonable time and resource constraints.
Validation of Gbsa for Screening Applications
Despite its limitations, GBSA has been extensively validated for screening applications in drug discovery and materials science.27 While absolute solvation energies may contain systematic errors, the relative ordering of molecules and structure-property trends are often preserved. This makes GBSA suitable for the primary objective of this study: identifying structural patterns and relative rankings rather than predicting absolute TADF efficiencies.
The key assumption underlying our approach is that the systematic errors introduced by the ground-state approximation affect all molecules similarly, preserving the relative trends essential for deriving design guidelines. This assumption is supported by the observation that GBSA models provide reliable relative free energies for structurally similar compounds.27
Limitations and Future Directions
We acknowledge several limitations of the GBSA approach for TADF emitters:
  1. 1.
    Quantitative accuracy: Absolute $\Delta E_{\text{ST}}$  values may deviate from experimental measurements due to inadequate treatment of CT state solvation.
  2. 2.
    Solvent dependence: The model cannot capture how different solvents might alter the relative rankings of molecules.
  3. 3.
    Strong CT states: Molecules with highly polarized CT states may be most affected by the ground-state approximation.
Future studies should validate these findings using state-specific solvation models for representative subsets of molecules, particularly those identified as promising TADF candidates. Additionally, experimental validation in different solvents would help assess the transferability of the derived design guidelines.
Justification for Current Approach
Given the screening nature of this study and the computational constraints, GBSA represents an appropriate balance between accuracy and efficiency. The model enables the systematic analysis of 747 molecules while maintaining sufficient accuracy to identify meaningful structure-property relationships. The derived guidelines should be viewed as screening criteria for identifying promising candidates, with the understanding that quantitative predictions require more sophisticated solvation treatments.
Specific Limitations for Multi-resonance Emitters
MR-TADF emitters, characterized by their extended $\pi$-conjugation and multiple resonance structures involving boron-nitrogen doping, are known to exhibit multi-reference character in their electronic structure.28, 29 While single-reference methods like sTD-DFT-xTB may not fully capture the quantitative energetics of these systems, previous validation studies have demonstrated that qualitative trends in $\Delta E_{\text{ST}}$  and orbital properties remain consistent between semiempirical, TD-DFT, and post-Hartree-Fock methods.30, 31
[…]
Spin-orbit Coupling Calculations
[…]
Derivation of Structure-property Descriptors and Theoretical Framework
@@
To translate the large computational data set into actionable design rules, donor-acceptor (D-A) torsional angles were automatically extracted from all optimized geometries. To quantify the degree of charge-transfer and its relation to geometry, we calculated the HOMO-LUMO overlap integral ($S^{\prime}_{\text{HL}}$) and the distance between the centroids of the hole and electron densities ($\Delta r_{\text{CT}}$) using the Multiwfn package 38, 39. Principal Component Analysis (PCA) was performed on the standardized matrix of all descriptors to identify the underlying variables governing TADF performance.
@@
This analysis is grounded in the theoretical framework of TADF efficiency, which depends not only on a small $\Delta E_{\text{ST}}$  (thermodynamics) but also on a high rate of reverse intersystem crossing ($k_{\text{RISC}}$), a kinetic parameter governed by spin-orbit coupling (SOC). According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation) 40, 41. Our chosen descriptors—particularly the D-A torsional angle—are direct proxies for tuning this CT/LE character balance. While our protocol does not compute SOC explicitly due to its prohibitive cost, analyzing these structural and electronic descriptors provides a robust, qualitative assessment of the kinetic factors that, alongside $\Delta E_{\text{ST}}$, govern overall TADF performance.
@@
RESULTS AND DISCUSSION: FROM STATISTICAL TRENDS TO RATIONAL DESIGN PRINCIPLES
@@
This section leverages the large-scale data set of $747$ TADF emitters, processed with the validated HTVS protocol detailed in our companion paper (Article 1 13), to extract quantitative, statistically robust structure-property relationships. Our analysis transitions from broad statistical trends to specific, actionable molecular design rules, with a focus on providing a deep mechanistic understanding that connects molecular structure to the key photophysical parameters governing TADF efficiency.
Low-dimensional Nature of the Tadf Design Space
@@
Architectural Hierarchy Defines Tadf Potential
@@
Data-driven Clustering Reveals Distinct Molecular Families
To complement the structure-based classification, we performed an unsupervised K-means clustering based on the molecules’ photophysical properties. This data-driven approach identifies four distinct molecular families, as shown in Figure˜3:
@@
Multi-resonance Tadf: a Distinct Paradigm for Blue Emitters
@@
However, as noted in the methodology, it is important to interpret these results with caution. The inherent electronic complexity of MR-TADF systems makes them prone to multi-reference character 42, which is not fully captured by the single-reference nature of the semiempirical methods used in this HTS. These results guide compound ranking but require higher-level validation for quantitative accuracy.
@@
Deconstructing Performance: Geometric and Electronic Design Levers
Torsional Angle as a Primary Control Parameter
@@
Quantifying the Torsion-induced Electronic Trade-off
This geometric control is directly reflected in the electronic structure. As shown in Figure˜5, twisting the molecule reduces the HOMO-LUMO overlap ($S^{\prime}_{\text{HL}}$), which correlates strongly with a smaller $\Delta E_{\text{ST}}$  (Spearman $\rho=-0.45$). However, this comes at a cost: the oscillator strength, essential for a high radiative rate, is directly proportional to this overlap ($\rho=0.78$). This visualization makes the fundamental TADF design trade-off explicit: minimizing the energy gap via orbital separation inherently suppresses emission intensity. An optimal region for high-performance emitters lies where there is intermediate overlap ($0.150.30$) that balances these competing requirements.
To move beyond this qualitative picture and establish a more rigorous quantitative link, future work could employ methods such as Natural Transition Orbitals (NTOs) or the RespA procedure 43, 44. These tools would allow for a compact representation of the virtual excitations and rigorously connect the molecular geometry to the precise charge-transfer character (e.g., via centroid distance $\Delta r$) and the ensuing coupling efficiency 45.
@@
Implications for Reverse Intersystem Crossing ($k_{\text{risc}}$)
While a small $\Delta E_{\text{ST}}$  is the primary thermodynamic requirement for TADF, the overall efficiency is governed by the kinetics of RISC. While direct calculation of the $k_{\text{RISC}}$ rate is outside the scope of HTS, the derived structural principles must be explicitly connected to the kinetics of TADF. Within the Marcus theory framework, $k_{\text{RISC}}$ is proportional to the square of the spin-orbit coupling (SOC) matrix element and exponentially dependent on $\Delta E_{\text{ST}}$  and the reorganization energy ($\lambda$) 46.
According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation, 1CT $\leftrightarrow$ 3LE) 40. This creates a mechanistic challenge: the very act of creating a pure CT state in $S_{1}$ to minimize $\Delta E_{\text{ST}}$  can lead to a $T_{1}$ state that also has pure CT character, making the transition spin-forbidden (1CT $\leftrightarrow$ 3CT) and suppressing $k_{\text{RISC}}$.
@@
HOMO-LUMO overlap Directly modulates $\Delta E_{\text{ST}}$ (thermodynamics) $0.150.30$
@@
Synthesis of a Data-driven Design Strategy
The combination of these large-scale statistical analyses allows us to formulate a set of quantitative, data-driven design guidelines for accelerating the discovery of novel TADF emitters, summarized in Table˜3. These rules provide a clear roadmap for molecular engineering, from selecting a high-performance architecture to fine-tuning its properties via geometric and electronic modifications.
@@
HOMO-LUMO Overlap ($S^{\prime}_{\text{HL}}$) $0.150.30$ Navigates the trade-off between $\Delta E_{\text{ST}}$  and oscillator strength.
@@
Future Outlook and Data Exploitation
The Hts Data Set as a Machine Learning Substrate
@@
Actionable Design Targets
@@
Sensitivity Analysis of Design Guidelines
To assess the robustness of our HTVS-derived design guidelines, we performed a comprehensive sensitivity analysis examining how variations in threshold choices affect the identification of promising TADF candidates. This analysis addresses the inherent subjectivity in defining quantitative cutoffs and provides confidence intervals for our screening criteria.
Threshold Sensitivity for $\Delta E_{\text{ST}}$  Criteria
We systematically varied the $\Delta E_{\text{ST}}$  threshold from 0.2 to 0.4 eV to examine its impact on candidate selection. The results demonstrate that threshold choice significantly influences the number of molecules classified as promising: $\Delta E_{\text{ST}}$  $\leq 0.2$ eV identifies 163 molecules (21.8%), while $\Delta E_{\text{ST}}$$\leq 0.4$ eV captures 507 molecules (67.9%). The 0.3 eV threshold employed throughout this study represents a balanced choice that captures 345 molecules (46.2%) while maintaining selectivity.
When combining $\Delta E_{\text{ST}}$  and HOMO-LUMO overlap criteria, the sensitivity becomes more pronounced. For example, applying both $\Delta E_{\text{ST}}$$\leq 0.3$ eV and overlap $\leq 0.4$ yields 272 molecules (36.4%), while the more restrictive combination identifies 134 molecules (17.9%). This demonstrates that our guidelines represent screening criteria rather than absolute predictors.